# From random walks to epidemic spreading: Compartment model with mortality for vector transmitted diseases [^1] PROCEEDINGS ICDEA 2024

### v1

@@ 3.1 Multiple walkers approach
In order to obtain a microscopic picture of the dynamics, we implement the mean field model into random walk simulations. We consider a population of $Z$ individuals, represented by random walkers moving independently on connected undirected (ergodic) networks [^3] Each pair of distinct nodes is connected by a path along edges of finite length. where we focus on random graphs of the Barabási–Albert (BA) and Watts– Strogatz (WS) type by using the PYTHON NetworkX library.
@@
The considered WS network has connectivity parameter $m=2$ with emergence of a large-world graph, see Figs. 3, 4. The WS graph is generated by starting in a first step by a ring of $N$ nodes, where each node is connected to its $m$ left and $m$ right closest neighbors, thus each node has $2m$ connections. In a second step, each of the connections is rewired with a probability $p$ to a randomly chosen node by avoiding multiple connections [17, 21]. The considered WS graph $G$ is generated by NetworkX employing $G=nx.connected\_watts\_strogatz\_graph(N,m,p,seed=seed)$ (allowing to fix the "seed", i.e. the generated random numbers to obtain the same random graph in different simulation runs).
@@
$K_{\alpha,\xi}(t)=\frac{\xi^{\alpha}t^{\alpha-1}}{\Gamma(\alpha)}e^{-\xi t},\hskip 28.45274pt\xi,\,\,\alpha>0$ (25)
where $\alpha$ is the so-called “shape parameter” and $\xi$ the rate parameter (often, the term “scale parameter” is used, $\theta=\xi^{-1}$) and $\Gamma(\alpha)$ stands for the Gamma function. The Gamma PDF has finite mean $\langle t\rangle_{\alpha,\xi}=\int_{0}^{\infty}tK_{\alpha,\xi}(t){\rm d}t=\frac{\alpha}{\xi}$, and for $\alpha\in(0,1)$, the Gamma PDF is weakly singular at $t=0$ and $\alpha=1$ recovers exponential PDFs. For $\alpha>1$, the Gamma PDF has a maximum at $t_{max}=\frac{\alpha-1}{\xi}$ and becomes narrower the larger $\xi$ approaching a Dirac $\delta$-distribution $K_{\alpha,\xi}(t)\to\delta(t-t_{0})$ for $\xi\to\infty$ while keeping its mean $\langle t\rangle_{\alpha,\xi}=t_{0}$ constant.
@@
In order to describe the random walk of each walker, we denote with $i=1,\ldots N$ the nodes of the network and introduce the $N\times N$ the symmetric adjacency matrix $(A_{ij})$ where $A_{ij}=1$ if the pair of nodes $i,j$ is connected by an edge, and $A_{ij}=0$ if the pair is disconnected. Further, we assume $A_{ii}=0$ to avoid self-connections of nodes. We restrict our analysis to undirected networks, where edges have no predefined direction and the adjacency matrix is symmetric. The degree $k_{i}$ of a node $i$ counts the number of its neighbor nodes (connected with $i$ by edges). Each walker performs independent Markovian steps between connected nodes. The steps from a node $i$ to one of its $k_{i}=\sum_{j=1}^{N}A_{ij}$ neighbor nodes are chosen with probability $1/k_{i}$, leading for all walkers to the same transition matrix which reads [19, 20, 32]
@@
Figure 6: Spreading without mortality in the same BA graph as in Fig. 5 ($1500$ walkers and averaging over $100$ equivalent RW realizations) and the same setting with identical parameters of Gamma distributed $t_{I}^{w,n}$ for some values of transmission probabilities $p_{n}$ of I-walkers to S-nodes (lower left and upper frames). This mimics the effect of confinement measures where $1-p_{n}$ is identified with the fraction of confined walkers’ population.
@@ 3.2 Discussion
We observe that the mean field model agrees well with the random walk result in the case of the small world BA graph (Figs. 5, 6). For the WS graph of Fig. 4 which is large world, the agreement between random walk and mean field is characterized by a delayed increase of the compartment fractions in the random walk simulations which can be explained that the infection needs more time to spread in a weakly connected architecture with long average distances between nodes. For large observation times, the agreement is well for zero mortality, but less well in presence of mortality. We explain this discrepancy that in the case of a large world graph the infection rates in the network are deviating from the simple mass-action-law forms (1).
@@ 4 Conclusions
Comparison of the results of the mean field model with random walk simulations shows that the mean field model assuming simple forms of the infection rates (Eqs. (1)) are well capturing the spreading dynamics in small world networks and vanishing mortality (see Figs. 5, 6). The agreement is less well in large world networks (Fig. 4) in the beginning of the epidemic, but again agrees well for large observation times in absence of mortality. Larger deviations between mean field and random walk occur for higher mortality. For future research, a more sophisticated mean field model would be desirable, which is able to better capture topological effects of large world (weakly connected) networks.

### v2

@@ 3.1 Multiple walkers approach
In order to obtain a microscopic picture of the dynamics, we implement the mean field model into random walk simulations. We consider a population of $Z$ individuals, represented by random walkers moving independently on connected undirected (ergodic) networks [^3] Each pair of distinct nodes is connected by a path along edges of finite length., where we focus on random graphs of the Barabási–Albert (BA) and Watts– Strogatz (WS) type by using the PYTHON NetworkX library.
@@
The considered WS network has connectivity parameter $m=2$ with emergence of a large-world graph, see Figs. 4, 5. The WS graph is generated by starting in a first step by a ring of $N$ nodes, where each node is connected to its $m$ left and $m$ right closest neighbors, thus each node has $2m$ connections. In a second step, each of the connections is rewired with a probability $p$ to a randomly chosen node by avoiding multiple connections [27, 28]. The considered WS graph $G$ is generated by NetworkX employing $G=nx.connected\_watts\_strogatz\_graph(N,m,p,seed=seed)$ (allowing to fix the "seed", i.e. the generated random numbers to obtain the same random graph in different simulation runs).
@@
$K_{\alpha,\xi}(t)=\frac{\xi^{\alpha}t^{\alpha-1}}{\Gamma(\alpha)}e^{-\xi t},\hskip 28.45274pt\xi,\,\,\alpha>0,$ (30)
where $\alpha$ is the so-called “shape parameter” and $\xi$ the rate parameter (often, the term “scale parameter” is used, $\theta=\xi^{-1}$) and $\Gamma(\alpha)$ stands for the Gamma function. The Gamma PDF has finite mean $\langle t\rangle_{\alpha,\xi}=\int_{0}^{\infty}tK_{\alpha,\xi}(t){\rm d}t=\frac{\alpha}{\xi}$, and for $\alpha\in(0,1)$, the Gamma PDF is weakly singular at $t=0$ and $\alpha=1$ recovers exponential PDFs. For $\alpha>1$, the Gamma PDF has a maximum at $t_{max}=\frac{\alpha-1}{\xi}$, and becomes narrower the larger $\xi$ approaching a Dirac $\delta$-distribution $K_{\alpha,\xi}(t)\to\delta(t-t_{0})$ for $\xi\to\infty$ while keeping its mean $\langle t\rangle_{\alpha,\xi}=t_{0}$ constant.
@@
In order to describe the random walk of each walker, we denote with $i=1,\ldots N$ the nodes of the network and introduce the $N\times N$ the symmetric adjacency matrix $(A_{ij})$, where $A_{ij}=1$ if the pair of nodes $i,j$ is connected by an edge, and $A_{ij}=0$ if the pair is disconnected. Further, we assume $A_{ii}=0$ to avoid self-connections of nodes. We restrict our analysis to undirected networks, where edges have no predefined direction and the adjacency matrix is symmetric. The degree $k_{i}$ of a node $i$ counts the number of its neighbor nodes (connected with $i$ by edges). Each walker performs independent Markovian steps between connected nodes. The steps from a node $i$ to one of its $k_{i}=\sum_{j=1}^{N}A_{ij}$ neighbor nodes are chosen with probability $1/k_{i}$, leading for all walkers to the same transition matrix which reads [15, 31, 32]
@@
Figure 7: Spreading without mortality in the same BA graph as in Fig. 6 ($1500$ walkers and averaging over $100$ equivalent RW realizations) and the same setting with identical parameters of Gamma distributed $t_{I}^{w,n}$ for some values of transmission probabilities $p_{n}$ of I-walkers to S-nodes (lower left and upper frames). This mimics the effect of confinement measures, where $1-p_{n}$ is identified with the fraction of confined walkers’ population.
@@ 3.2 Discussion
We observe that the mean field model agrees well with the random walk result in the case of the small world BA graph (Figs. 6, 7). For the WS graph of Fig. 5 which is large world, the agreement between random walk and mean field is characterized by a delayed increase of the compartment fractions in the random walk simulations which can be explained that the infection needs more time to spread in a weakly connected architecture with long average distances between nodes. For large observation times, the agreement is well for zero mortality, but less well in presence of mortality. We explain this discrepancy that in the case of a large world graph the infection rates in the network are deviating from the simple mass-action-law forms (1).
The validity of mass-action laws for the infection rates requires that the compartmental populations are fully mixed and spatially homogeneous. The deviation from mass-action law can be explained by the observation that in large world graphs it needs a certain delay time until the compartmental populations are sufficiently well mixed such that the local infection rates become homogeneous in the whole network and independent from position nodes of the walkers. Therefore, the infection rates take mass-action laws asymptotically in the large-time limit. This retardation effect can be clearly identified in Fig. 5, see especially left frames without mortality.
@@ 4 Conclusions
Comparison of the results of the mean field model with random walk simulations shows that the mean field model assuming simple forms of the infection rates (Eqs. (1)) are well capturing the spreading dynamics in small world networks and vanishing mortality (see Figs. 6, 7). The agreement is less well in large world networks (Fig. 5) in the beginning of the epidemic, but again agrees well for large observation times in absence of mortality. Larger deviations between mean field and random walk occur for higher mortality. For future research, a more sophisticated mean field model would be desirable, which is able to better capture topological effects of large world (weakly connected) networks. A promising direction to advance with this task is to extend our model to a degree-based mean-field approach [34]. In this way topological information of the network is incorporated which might avoid the drawbacks of mass-action law infection rates for small observation times.